Please help me solve this problem.
The measure of an angle is 3/4 the measure od its complement. Find the measure of the angle and its complement.

Answer by madhujiithcharu@yahoo.co.in(28)

Let A be the angle. Therefore A = 3/4(180-A)
A = 3x45-3/4A
A+3/4A = 135
7/4A = 135
A = 135 x 4/7
A = 540/7
A = 77 approximately and its complement is 103 approx.